# Simulation of Hypergraph Algorithms with Looped Transformers

### v1

@@
Simulation of Hypergraph Algorithms with Looped Transformers
Xiaoyu Li 7.xiaoyu.li@gmail.com. Independent Researcher.    Yingyu Liang yingyul@hku.hk. The University of Hong Kong. yliang@cs.wisc.edu. University of Wisconsin-Madison.    Jiangxuan Long lungchianghsuan@gmail.com. South China University of Technology.    Zhenmei Shi zhmeishi@cs.wisc.edu. University of Wisconsin-Madison.    Zhao Song magic.linuxkde@gmail.com. The Simons Institute for the Theory of Computing at the University of California, Berkeley.    Zhen Zhuang zhuan244@umn.edu. University of Minnesota.
Looped Transformers have shown exceptional capability in simulating traditional graph algorithms, but their application to more complex structures like hypergraphs remains underexplored. Hypergraphs generalize graphs by modeling higher-order relationships among multiple entities, enabling richer representations but introducing significant computational challenges. In this work, we extend the Loop Transformer architecture to simulate hypergraph algorithms efficiently, addressing the gap between neural networks and combinatorial optimization over hypergraphs. In this paper, we extend the Loop Transformer architecture to simulate hypergraph algorithms efficiently, addressing the gap between neural networks and combinatorial optimization over hypergraphs. Specifically, we propose a novel degradation mechanism for reducing hypergraphs to graph representations, enabling the simulation of graph-based algorithms, such as Dijkstra’s shortest path. Furthermore, we introduce a hyperedge-aware encoding scheme to simulate hypergraph-specific algorithms, exemplified by Helly’s algorithm. The paper establishes theoretical guarantees for these simulations, demonstrating the feasibility of processing high-dimensional and combinatorial data using Loop Transformers. This work highlights the potential of Transformers as general-purpose algorithmic solvers for structured data.
@@ 1 Introduction
Large Language Models (LLMs) built on the Transformer architecture [61], such as GPT-4o [47], Claude [1], and OpenAI’s latest o1 and o3 [48], demonstrate the remarkable potential to improve our daily lives. They are transforming domains like AI agents [12], search AI [48], AI assistants [46, 74], and conversational AI [31]. As their capabilities expand, LLMs serve as powerful instruments driving the future of technological innovation. A key emergent capability of LLMs is their adeptness at executing traditional algorithms, showcasing exceptional performance in tasks such as sorting large datasets and solving intricate problems. LLMs have shown particular aptitude in processing diverse data structures and types, including graph-structured data [69], tabular formats [67], and temporal sequences [78]. Notably, the Loop Transformer architecture has demonstrated robust performance in simulating various graph algorithms [13], achieving comparable results to traditional graph processing methods. These developments not only advance our theoretical understanding of Transformer-based models’ computational capabilities but also offer practical alternatives for complex algorithmic tasks. This success raises a compelling question:
Can Loop Transformers extend their algorithmic simulation capabilities to the more complex domain like hypergraphs?
Hypergraphs, which extend traditional graphs by permitting edges, called hyperedges, to connect multiple nodes, offer a more natural representation of many real-world datasets and problems. This added complexity enables hypergraphs to capture higher-order relationships between entities, making them particularly useful in a variety of domains where interactions are not limited to pairwise connections. For instance, in combinatorial optimization, hypergraphs are used to model complex systems such as resource allocation [76] and network flow [72], where multiple entities interact simultaneously in a single operation. Similarly, in recommendation systems [68, 66], hypergraphs can represent user-item interactions more flexibly, enabling the modeling of complex, multi-way relationships, such as collaborative filtering or group-based recommendations. In computational biology, hypergraphs are employed to capture intricate biological relationships, such as protein-protein interaction networks or gene expression data [60], where multiple factors simultaneously contribute to the system’s behavior.
Hypergraph algorithms, however, often pose greater computational challenges than graph algorithms due to their richer structure and exponential growth in possible hyperedge combinations. Bridging the gap between hypergraph algorithms and neural network simulation could unlock new paradigms for solving high-dimensional and combinatorially complex problems.
1.1 Our Contribution
In this paper, we extend the results of [13] that demonstrated the ability of Loop Transformers to simulate graph algorithms. We propose and investigate the hypothesis that Loop Transformers, with their iterative processing and representational flexibility, are capable of simulating hypergraph algorithms. By leveraging the inherent capacity of Transformers for parallel processing and multi-head attention, we aim to show that these models can effectively encode and operate over hypergraph structures. In detail, we show that the Loop Transformer can simulate a degradation mechanism (see detail in Section 6) by using 10 layers, where each attention layer includes three attention heads with constant feature dimensions. Also, the Loop Transformer can represent the Helly algorithm (see detail in Section 6) by using 11 attention layers, where each layer includes three attention heads with constant feature dimensions. These findings demonstrate that the Looped Transformer is a general-purpose computational tool capable of bridging neural networks and combinatorial optimization tasks on structured, high-dimensional data.
The contributions of our work are outlined as follows:
  • We design a degradation mechanism (Theorem 6.1) to simulate graph-based algorithms, e.g., Dijkstra, BFS, DFS, on hypergraphs via reduced graph representations within the Loop Transformer framework.
@@ 1.1 Our Contribution
  • We propose a novel encoding scheme to simulate hyperedge-aware algorithms, e.g., Helly’s algorithm (Theorem 6.2), by incorporating hyperedge-specific operations into Transformer’s iterative processes.
@@
We introduce the works related to our paper in Section 2. In Section 3, we present the basic concepts, e.g., hypergraph and Loop Transformers. In Section 4, we introduce the construction of some basic operations. In Section 5, we present some simulation results based on the basic operation. In Section 6, we detail our main results. In Section 7, we give a conclusion of our paper.
@@ 2.1 Neural Network Execute Algorithms
In recent years, the integration of neural networks with algorithmic execution has gained significant attention, leading to the emergence of a body of work aimed at improving the efficiency and performance of algorithmic tasks through deep learning techniques. This line of research focuses on leveraging the power of neural networks to either approximate traditional algorithms or directly replace them with learned models. [53] established the computational universality of finite recurrent neural networks with sigmoidal activation functions by proving their capability to simulate any Turing Machine. Later, [49] proved that Transformer architectures achieve Turing completeness through hard attention mechanisms, which enable effective computation and access to internal dense representations. Building with the framework of looped transformer, [18] uses looped transformers as the building block to make a programmable computer, showcasing the latent capabilities of Transformer-based neural networks. [13] demonstrated that looped transformers can implement various graph algorithms, including Dijkstra’s shortest path, Breadth-First Search (BFS), Depth-First Search (DFS), and Kosaraju’s algorithm for strongly connected components, through their multi-head attention mechanism. [18, 70] uses Transformers as the building block to build a programmable computer, showcasing the latent capabilities of Transformer-based neural networks. Beyond [18], [41] proves that a looped 23-layer $\mathrm{ReLU-MLP}$ is capable of performing the basic necessary operation of a programmable computer. [19, 9] showed that looped transformers can efficiently do in-context learning by multi-step gradient descent.
@@ 2.2 Hypergraphs in Neural Networks
Hypergraphs have recently emerged as a powerful mathematical model for representing complex relationships in data, and they have found promising applications in deep learning. One of the key motivations for using hypergraphs in neural networks is their ability to capture higher-order interactions among entities, which can be more representative of real-world data than traditional pairwise relationships. Several works have explored leveraging hypergraph-based representations to improve the performance of neural architectures in various domains. For instance, [16] introduced a hypergraph-based approach to enhance graph neural networks (GNNs), where the nodes of a graph were represented as hypernodes in a hypergraph. This structure allowed the model to better capture multi-way interactions, significantly improving results on tasks such as node classification and link prediction. In another study, [71] extended convolutional neural networks (CNNs) to work on hypergraphs, demonstrating how hypergraph convolution can be used to aggregate features over hyperedges, yielding more accurate feature representations in image and video processing tasks. Additionally, [6] proposed a novel method for integrating hypergraph attention mechanisms into deep neural networks, allowing the model to focus on the most influential hyperedges for decision-making. This approach was shown to enhance the interpretability of deep learning models and achieve state-of-the-art performance in tasks involving large-scale multi-modal data. In [4], the authors investigated the use of hypergraphs in the context of deep reinforcement learning, where the structure of the hypergraph facilitated more efficient exploration of state-action spaces, leading to faster convergence in complex environments. Recent relevant research includes works on tensor transformers [51, 3, 44, 33, 75], acceleration techniques [64, 63, 56, 36, 50, 59, 22, 57, 52, 62, 29, 23, 24, 35, 27, 7, 30, 37, 34, 26, 42, 40, 28, 39, 2, 11], and other related studies [14, 54, 20, 43, 58, 8, 65, 45, 15, 17, 10, 32, 25, 55, 77, 73, 38].
@@ 3.2 Simulation
Definition 3.1 (Simulation, Definition 3.1 on page 3 in [13]).
@@ 3.4 Looped Transformers
Definition 3.4 (Single-head attention, [13]).
@@
Definition 3.6 (Multi-head attention, [13]).
@@
$f_{\mathrm{attn}}(X,\widetilde{A})~{}:=~{}\sum_{i\in M_{A}}\psi^{(i)}(X,\widetilde{A})+\sum_{i\in M_{A^{\top}}}\psi^{(i)}(X,\widetilde{A}^{\top})+\sum_{i\in M}\psi^{(i)}(X,I_{n+1})+X,$
@@
Remark 3.10.
In the construction of this paper, we set $X\in\mathbb{R}^{K\times d}$, where $K$ is defined in Definition 3.3, $d$ is a constant independent of $K$. The matrix $X$ stores different variables in its different columns. We use $B_{\mathrm{global}}$ to present the column that only the top one scalar is $1$ while the rest are 00. We use $B_{\mathrm{local}}$ to present the column that only the top one scalar is 00 while the rest are $1$. We use $P$ to present the 2 columns of position embedding, where the first column is $\sin(\theta_{i})$, and the second column is $\cos(\theta_{i})$ for $i\in[K-1]$. We use $P_{\mathrm{cur}}$ to present the position embedding of the current vertex or hyperedge.
@@
Definition 3.11 (Transformer layer, [13]).
@@
Definition 3.12 (Multi layer Transformer, [13]).
Let $\widetilde{A}$ be defined in Definition 3.3, let $m=O(1)$ denote the layer of transformers. Let the transformer layer $f$ be defined in 3.11. We define the multi-layer transformer as
@@
Algorithm 2 Iteration of minimum value
3:$\mathrm{visit}_{\min}\leftarrow{\bf 0}_{n_{v}}\|{\bf 1}_{K-1-n_{v}}$
5:
6:procedure GetMinimumValue($x\in\mathbb{R}^{d}$)
7:     if $\mathrm{termination}_{\min}=1$ then
8:         $\cdots$ $\triangleright$ Re-initialization of minimum value, Lemma 4.1
9:         $\mathrm{termination}_{\min}\leftarrow 0$
10:     end if
11:     $\mathrm{idx}_{\mathrm{cur}}\leftarrow\mathrm{idx}_{\mathrm{cur}}+1$triangleright Increment, Lemma 4.2
@@
12:     $\mathrm{val}_{\mathrm{cur}}\leftarrow x[\mathrm{idx}_{\mathrm{cur}}]$triangleright Read scalar from column, Lemma 4.3
13:     if $\mathrm{val}_{\mathrm{cur}}<\mathrm{val}_{\mathrm{best}}$ $\triangleright$ Compare, Lemma 4.4
@@
14:         $\mathrm{val}_{\mathrm{best}},~{}\mathrm{idx}_{\mathrm{best}}\leftarrow\mathrm{val}_{\mathrm{cur}},~{}\mathrm{idx}_{\mathrm{cur}}$triangleright Update variables, Lemma 4.1
15:     end if
@@
16:     $\mathrm{visit}_{\min}[\mathrm{idx}_{\mathrm{cur}}]\leftarrow 1$triangleright Write scalar to column, Lemma 4.5
@@
17:     $\mathrm{termination}_{\min}\leftarrow\neg(0~{}\mathrm{in}~{}\mathrm{visit}_{\min})$triangleright Trigger termination, Lemma 4.6
18:end procedure

### v2

# Neural Algorithmic Reasoning for Hypergraphs with Looped Transformers

@@
Neural Algorithmic Reasoning for Hypergraphs with Looped Transformers
Xiaoyu Li xiaoyu.li2@student.unsw.edu.au. University of New South Wales.    Yingyu Liang yingyul@hku.hk. The University of Hong Kong. yliang@cs.wisc.edu. University of Wisconsin-Madison.    Jiangxuan Long lungchianghsuan@gmail.com. South China University of Technology.    Zhenmei Shi zhmeishi@cs.wisc.edu. University of Wisconsin-Madison.    Zhao Song magic.linuxkde@gmail.com. The Simons Institute for the Theory of Computing at the University of California, Berkeley.    Zhen Zhuang zhuan244@umn.edu. University of Minnesota.
Looped Transformers have shown exceptional neural algorithmic reasoning capability in simulating traditional graph algorithms, but their application to more complex structures like hypergraphs remains underexplored. Hypergraphs generalize graphs by modeling higher-order relationships among multiple entities, enabling richer representations but introducing significant computational challenges. In this work, we extend the Loop Transformer architecture’s neural algorithmic reasoning capability to simulate hypergraph algorithms, addressing the gap between neural networks and combinatorial optimization over hypergraphs. Specifically, we propose a novel degradation mechanism for reducing hypergraphs to graph representations, enabling the simulation of graph-based algorithms, such as Dijkstra’s shortest path. Furthermore, we introduce a hyperedge-aware encoding scheme to simulate hypergraph-specific algorithms, exemplified by Helly’s algorithm. We establish theoretical guarantees for these simulations, demonstrating the feasibility of processing high-dimensional and combinatorial data using Loop Transformers. This work highlights the potential of Transformers as general-purpose algorithmic solvers for structured data.
@@ 1 Introduction
Algorithms underlie many of the technological innovations that shape our daily lives, from route planners [43] to search engines [9]. In parallel, the advent of Large Language Models (LLMs), which are built on the Transformer architecture [97], has dramatically extended the boundary of what is achievable through machine learning. Noteworthy systems such as GPT-4o [75], Claude [2], and the latest OpenAI models [77] have already propelled advances across a variety of domains, including AI agents [20], AI search [77], and conversational AI [72, 114, 54]. As LLMs become increasingly capable, they open up new possibilities for integrating algorithmic reasoning into neural models.
Building on this idea, neural algorithmic reasoning [94] has recently emerged as a means to fuse the abstraction and guarantees of algorithms with the representational flexibility of neural networks. By learning to mimic the step-by-step operations of well-known procedures, neural networks can capture the generalization capabilities of algorithms. Following a self-regression pattern, [23] introduces an embedding representation for iterative graph algorithms, facilitating scalable learning of steady-state solutions. Meanwhile, [49] employs reinforcement learning and graph embedding to automate heuristic design for NP-hard combinatorial optimization problems. By supervising the network on each intermediate state rather than only the end result, [98] shows how learned subroutines can transfer across related tasks like Bellman-Ford algorithm [6] and Prim’s algorithm [80]. Encouraging results have also been obtained in simulating graph algorithms using looped Transformers [24], suggesting that Transformers may be capable of executing procedural knowledge at scale.
Extending this paradigm to more complex relational structures, such as hypergraphs, presents new opportunities and challenges. Hypergraphs naturally allow each hyperedge to contain multiple vertices, making them especially suitable for capturing higher-order interactions present in many real-world tasks. Recent studies [104, 28, 110] illustrate how hypergraph-based representations can enhance relational reasoning, manage complex data interdependencies, and support a larger subset of relational algebra operations. However, while standard graphs often rely on symmetrical adjacency matrices to encode connections, hypergraphs are typically represented by incidence matrices, posing additional challenges. For instance, Helly’s algorithm [41], which assesses whether a hypergraph exhibits the Helly property [11, 95], requires operations over these incidence representations and has been shown to be NP-Hard in certain extensions [26]. Similarly, converting hypergraphs to graph-like structures often inflates feature dimensions linearly with the number of vertices due to matrix-storage demands [10, 55]. These insights collectively raise a central question:
Can looped Transformers achieve neural algorithmic reasoning on hypergraphs
using $O(1)$ feature dimensions and $O(1)$ layers?
Our Contribution.
In this paper, we answer this question affirmatively by extending the results in [24], which demonstrated the ability of Loop Transformers to simulate graph algorithms.
We propose and investigate the hypothesis that Loop Transformers, with their iterative processing and representational flexibility, are capable of simulating hypergraph algorithms. By leveraging the inherent capacity of Transformers for parallel processing and multi-head attention, we aim to demonstrate that these models can effectively encode and operate over hypergraph structures. The contributions of our work are outlined as follows:
  • We design a degradation mechanism (Theorem 4.1) to simulate graph-based algorithms, such as Dijkstra, BFS, and DFS, on hypergraphs by dynamically degrading the incidence matrix to an adjacency matrix implicitly. This process can be simulated by a looped Transformer with constant layers and constant feature dimensions.
@@
  • We propose a novel encoding scheme to simulate Helly’s algorithm (Theorem 4.2), which incorporates hyperedge-specific operations into the iterative processes of Transformers. This can also be simulated by a looped Transformer with constant layers and constant feature dimensions.
@@
We introduce the works related to our paper in Section 2. In Section 3, we present the basic concepts, e.g., hypergraph and Loop Transformers. In Section 4, we detail our main results. In Section 5, we introduce the construction of some basic operations. In Section 6, we present some simulation results based on the basic operation. In Section 7, we give a conclusion of our paper.
@@ 2.1 Neural Network Execute Algorithms
In recent years, the integration of neural networks with algorithmic execution has gained significant attention, leading to the emergence of a body of work aimed at improving the efficiency and performance of algorithmic tasks through deep learning techniques. This line of research focuses on leveraging the power of neural networks to either approximate traditional algorithms or directly replace them with learned models. [85] established the computational universality of finite recurrent neural networks with sigmoidal activation functions by proving their capability to simulate any Turing Machine. Later, [79] proved that Transformer architectures achieve Turing completeness through hard attention mechanisms, which enable effective computation and access to internal dense representations. Building with the framework of looped transformer, [34] uses looped transformers as the building block to make a programmable computer, showcasing the latent capabilities of Transformer-based neural networks. [24] demonstrated that looped transformers can implement various graph algorithms, including Dijkstra’s shortest path, Breadth-First Search (BFS), Depth-First Search (DFS), and Kosaraju’s algorithm for strongly connected components, through their multi-head attention mechanism.
@@ 2.2 Hypergraphs in Neural Networks
Hypergraphs have recently emerged as a powerful mathematical model for representing complex relationships in data, and they have found promising applications in deep learning. Several works have explored leveraging hypergraph-based representations to improve the performance of neural architectures in various domains. For instance, [29] introduced a hypergraph-based approach to enhance graph neural networks (GNNs), enabling better multi-way interaction modeling for tasks like node classification and link prediction, while subsequent works extended convolutional networks to hypergraphs [109], incorporated hypergraph attention for improved interpretability [12], and leveraged hypergraphs in deep reinforcement learning for faster convergence [7]. Hypergraphs have also been employed as a powerful framework for advanced reasoning. Recent research [104, 28, 110] elucidates their capacity to refine relational inference, encapsulate intricate data interdependencies, and facilitate a broader spectrum of operations within relational algebra, thereby augmenting the expressiveness and computational efficacy of reasoning paradigms.
2.3 Looped Transformer
First introduced by [22], the Universal Transformer, a variant of the Transformer with a recurrent inductive bias, can be regarded as a foundational model for looped Transformers. Empirical evidence from [108] suggests that increasing the number of loop iterations enhances performance in various data-fitting tasks while requiring fewer parameters. Furthermore, [35] establishes a theoretical guarantee that the looped Transformer can execute multi-step gradient descent, thereby ensuring convergence to algorithmic solutions. Recent research [1, 37] has deepened our understanding of how specific algorithms can be emulated and how their training dynamics facilitate convergence, particularly in the context of in-context learning. [35, 16] showed that looped transformers can efficiently do in-context learning by multi-step gradient descent. [34] uses Transformers as the building block to build a programmable computer, showcasing the latent capabilities of Transformer-based neural networks. Beyond [34], [66] proves that a looped 23-layer $\mathrm{ReLU-MLP}$ is capable of performing the basic necessary operation of a programmable computer.
@@ 3.2 Simulation
Definition 3.1 (Simulation, Definition 3.1 in [24]).
@@ 3.4 Looped Transformers
Definition 3.4 (Single-head attention).
@@
Definition 3.6 (Multi-head attention).
@@
$f_{\mathrm{attn}}(X,\widetilde{A})~{}$ $:=~{}\sum_{i\in M_{A}}\psi^{(i)}(X,\widetilde{A})+\sum_{i\in M_{A^{\top}}}\psi^{(i)}(X,\widetilde{A}^{\top})$
$+\sum_{i\in M}\psi^{(i)}(X,I_{n+1})+X,$
@@
Definition 3.10 (Transformer layer).
@@
Definition 3.11 (Multi layer Transformer).
Let $\widetilde{A}$ be defined in Definition 3.3, let $m=O(1)$ denote the layer of transformers. Let the transformer layer $f$ be defined in 3.10. We define the multi-layer transformer as
@@
In the construction of this paper, we set $X\in\mathbb{R}^{K\times d}$, where $K$ is defined in Definition 3.3, and $d$ is a constant independent of $K$. The matrix $X$ stores different variables in its columns. A variable can either be an array or a scalar. Scalars are stored in the top row of the corresponding column, leaving the remaining $K-1$ rows as 00. Arrays are stored in the bottom $K-1$ rows of the column, leaving the top row as 00. We use $B_{\mathrm{global}}$ to represent a column where only the top scalar is $1$, while the rest of the entries are 00. Conversely, we use $B_{\mathrm{local}}$ to represent a column where the top scalar is 00, while the remaining entries are $1$. Additionally, we use $P$ to represent the two columns of position embeddings, where the first column contains $\sin(\theta_{i})$, and the second column contains $\cos(\theta_{i})$ for $i\in[K-1]$. Finally, $P_{\mathrm{cur}}$ is used to represent the position embedding of the current vertex or hyperedge.
@@
4 Main Results
4.1 Degradation
We observe that [24] presents the running results of several algorithms on graphs, which primarily include Dijkstra’s algorithm for the shortest path, Breadth-First Search (BFS), and Depth-First Search (DFS). [24] provides a general representation of such algorithms in Algorithm 7 of their paper. Given that Dijkstra’s algorithm for the shortest path, BFS, and DFS can be straightforwardly extended to hypergraphs [38] by identifying the shortest hyperedge between two nodes and treating it as the distance between those nodes, we propose a degradation mechanism in Algorithm 2. This mechanism allows dynamic access to the shortest hyperedge between two vertices without storing static adjacent matrix. Using this mechanism, we can extend Dijkstra’s algorithm for the shortest path, BFS, and DFS from graphs to hypergraphs. We formally state the theorem regarding the degradation mechanism as follows:
Algorithm 2 Degradation Iterate of hyperedge
1:$\mathrm{idx}_{\mathrm{hyperedge}}\leftarrow 0$
2:$\mathrm{val}_{\mathrm{hyperedge}}\leftarrow{\bf 0}_{n_{v}}\|{\bf 1}_{K-1-n_{v}}$
3:$\mathrm{iszero}_{\mathrm{hyperedge}}\leftarrow{\bf 0}_{n_{v}}\|{\bf 1}_{K-1-n_{v}}$
4:$\mathrm{candidates}_{\mathrm{hyperedge}}\leftarrow\Omega\cdot{\bf 1}_{K-1}$
5:$\mathrm{visit}_{\mathrm{hyperedge}}\leftarrow{\bf 0}_{n_{e}}\|{\bf 1}_{K-1-n_{e}}$
6:$\mathrm{termination}_{\mathrm{hyperedge}}\leftarrow 0$triangleright Initialization of visiting hyperedges
7:
8:procedure VisitHyperedge($\widetilde{A}\in\mathbb{R}_{+}^{n_{v}\times n_{e}}$)
9:     if $\mathrm{termination}_{\mathrm{hyperedge}}=1$ then
10:         $\cdots$ $\triangleright$ Re-initialization of visiting edge, Lemma C.1
11:         $\mathrm{termination}_{\min}\leftarrow 0$
12:     end if
13:     $\mathrm{idx}_{\mathrm{hyperedge}}\leftarrow\mathrm{idx}_{\mathrm{hyperedge}}+1$triangleright Increment, Lemma C.2
@@ 4.1 Degradation
14:     $\mathrm{val}_{\mathrm{hyperedge}}\leftarrow A[:,\mathrm{idx}_{\mathrm{hyperedge}}]$triangleright Read column from incident matrix, Lemma C.7
15:     for $i=1\to K-1$ do
16:         $\mathrm{iszero}_{\mathrm{hyperedge}}[i]\leftarrow(\mathrm{val}_{\mathrm{hyperedge}}[i]\leq 0)$triangleright Compare, Lemma C.3
17:     end for
18:     for $i=1\to K-1$ do
19:         if $\mathrm{iszero}_{\mathrm{hyperedge}}[i]=1$  then
20:              $\mathrm{val}_{\mathrm{hyperedge}}[i]\leftarrow\Omega$triangleright Selection, Lemma C.1
21:         end if
22:     end for
23:     for $i=1\to K-1$ do
24:         if $\mathrm{val}_{\mathrm{hyperedge}}[i]<\mathrm{candidates}[i]$ then $\triangleright$ Compare, Lemma C.3
@@
25:              $\mathrm{candidates}[i]\leftarrow\mathrm{val}_{\mathrm{hyperedge}}[i]$triangleright Update variables, Lemma C.1
26:         end if
27:     end for
@@
28:     $\mathrm{visit}_{\mathrm{hyperedge}}[\mathrm{idx}_{\mathrm{hyperedge}}]\leftarrow 1$triangleright Write scalar to column, Lemma C.5
@@
29:     $\mathrm{termination}_{\mathrm{hyperedge}}\leftarrow\neg(0~{}\mathrm{in}~{}\mathrm{visit}_{\min})$triangleright Trigger termination, Lemma C.6
30:     $\mathrm{termination}_{\min}\leftarrow\mathrm{termination}_{\min}\land\mathrm{termination}_{\mathrm{hyperedge}}$triangleright AND, Lemma C.8
31:end procedure
Theorem 4.1 (Degradation, informal version of Theorem E.1).
A looped transformer $h_{T}$ defined in Definition 3.11 exists, consisting of 10 layers, where each layer includes 3 attention heads with feature dimension of $O(1)$. This transformer can simulate the degradation operation (Algorithm 2) for hypergraphs, supporting up to $O(\widehat{\delta}^{-1})$ vertices and $O(\widehat{\delta}^{-1})$ hyperedges, where $\widehat{\delta}$ defined in Definition 3.12 is the nearest representable approximation of the minimum increment angle.
4.2 Helly
The Helly property in a hypergraph refers to a specific condition in the family of its hyperedges. A hypergraph is said to satisfy the Helly property if, for every collection of its hyperedges, whenever the intersection of every pair of hyperedges in the collection is non-empty, there exists at least one hyperedge in the collection that intersects all the others. This property is named after Helly’s theorem in convex geometry, which inspired its application in combinatorial settings like hypergraphs. We have reformulated the algorithm from Algorithm 3 of [10], which determines whether a hypergraph possesses the Helly property, into a form executable by the Looped Transformer, represented as our Algorithm 5. We now state the following theorem:
Theorem 4.2 (Helly, informal version of Theorem E.2).
A looped transformer $h_{T}$ exists, where each layer is defined as in Definition 3.10, consisting of 11 layers, where each layer includes 3 attention heads with feature dimension of $O(1)$. This transformer can simulate the Helly algorithm (Algorithm 5) for weighted hypergraphs, handling up to $O(\widehat{\delta}^{-1})$ vertices and $O(\widehat{\delta}^{-1})$ hyperedges, where $\widehat{\delta}$ defined in Definition 3.12 is the nearest representable approximation of the minimum increment angle.
4.3 Furthure Discussion for the Power of Looped Transformer
Our algorithm can be extended to the family of deterministic hypergraph motifs algorithms introduced by [55]. By using a similar proof as in Theorem 4.2, we can simulate Hypergraph Projection as a preprocessing step to obtain a static projection graph and simulate the Exact H-motif Counting operation based on this static projection graph. Combining these observations, we conclude that our algorithm can effectively simulate deterministic hypergraph motifs algorithms. Specifically, Hypergraph Projection can be simulated using the selection operation in Lemma C.1 and the addition operation in Lemma C.10; Exact H-motif Counting can be simulated using the AND operation in Lemma C.8, the comparison operation in Lemma C.3, and the addition operation in Lemma C.10. Note that although our method can leverage a looped Transformer to simulate deterministic algorithms on hypergraphs, it cannot simulate algorithms involving stochastic processes or sampling, such as random sampling on hypergraphs (see Algorithm 3 in [55]).
Furthermore, we can dynamically compute the hypergraph projection in a looped process, as presented in Theorem 4.1. The key difference lies in the computational resources: the static projection graph requires $O(K)$ columns for storage due to its size dependence on $K$, which prevents a solution in $O(1)$ column space. However, if we employ a similar method to Algorithm 2, the storage cost reduces to $O(1)$ columns at the expense of requiring more iterations. Thus, our methods may cover all sets of deterministic hypergraph motifs and show that the looped transfer is powerful.